OpenStudy (anonymous):

whats 19 over 57 equals? fractions--

OpenStudy (anonymous):

19 /57 0.333

jimthompson5910 (jim_thompson5910):

notice how 57 = 19*3 So this means that \[\Large \frac{19}{57}\] \[\Large \frac{19}{19*3}\] \[\Large \frac{\cancel{19}}{\cancel{19}*3}\] \[\Large \frac{1}{3}\] This means that \[\Large \frac{19}{57}=\frac{1}{3}\]